OpenStudy (anonymous):

4y-8x+12=0 find the slope and y-intercept

OpenStudy (anonymous):

You find the y intercept easily by putting x=0 and solving for y.

OpenStudy (anonymous):

4y=8x+12 y=2x+3 Therefore y intercept is 12 and slope is 2

OpenStudy (anonymous):

yintercept is wrong

OpenStudy (anonymous):

Yes, the y intercept is (0,3), since 4y=8*(0)+12 -> y=3

OpenStudy (anonymous):

-3

OpenStudy (id21):

\[4y-8x+12=0\] \[4y=8x-12\] y=2x-3 slope=2 y intercept=-3

OpenStudy (anonymous):

id21 is correct of course, sorry for the mistake!